OpenStudy (anonymous):

A motorist drives north for 34.0 minutes at 61.5 km/h and then stops for 15.0 minutes. He then continues north, traveling 130 km in 2.20 h. What is his average velocity? The answer is 54.6 but i cant seem to get it

OpenStudy (anonymous):

are you converting the time into the same units?

OpenStudy (anonymous):

\[\frac{d}{t_{total}}=V_{avg} \]What is the distance he travelled in his first leg? \[V=^d/_t\]\[61.5^{km}/_h=\frac{d}{.567h}\]You get 34.9 km. He travelled no distance in 15 minutes. This is going to lower is average velocity. The last leg does the work for you that we did in the first leg; we know the distance and the time!\[\frac{34.9+130}{34+15+132}\]You get .911 kilometers per minute. Multiply by 60kph to get 54.7kph! Hope this helps.

OpenStudy (anonymous):

.567 hours is 34 minutes divided by 60 minutes/hour btw

OpenStudy (anonymous):

d t total =V avg What is the distance he travelled in his first leg? V= d / t 61.5 km / h =d .567h You get 34.9 km. He travelled no distance in 15 minutes. This is going to lower is average velocity. The last leg does the work for you that we did in the first leg; we know the distance and the time!34.9+130 34+15+132 You get .911 kilometers per minute. Multiply by 60kph to get 54.7kph fred33

OpenStudy (anonymous):

AverageSpeed=Total Distance Driven divided by the Total Time consumed.\[\text{AverageSpeed}=\frac{\left(61.5*\frac{34}{60}\right)+130}{\frac{34}{60}+\frac{15}{60}+\frac{132}{60}}=54.6464 \text{kph} \]